Let's assume there are no taxes. Suppose the WACC is 12%. The debt ratio (i.e., weight on debt) is 60%. The interest rate is 7%. What is the expected return on equity?
Answer #1

Weight of equity=100-60=40%

WACC=Respective weight*Respective return

12=(0.6*7)+(0.4*Return on equity)

Return on equity=(12-4.2)/0.4

=19.5%